A spring obeying Hooke's law has a force constant k.Now if th espring is cut into two equal parts ,find the force constant of each part

Grade:11

1 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Rahul,

 

If spring is cut inti M:N then

K1  = k * (M+N/M) and K2 = k * (M+N/N)

as K is inversly proportional to length.

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay